OpenStudy (anonymous):

Help? Simplify 8x^3-6=119

OpenStudy (anonymous):

What a common factor between 6 and 8?

OpenStudy (shamim):

plz c my attachment

OpenStudy (shamim):

feel free to ask if any confusion

OpenStudy (anonymous):

common factor is 2 right?

OpenStudy (shamim):

did u c my solution?

OpenStudy (anonymous):

ohh yes thank you:) i have another question.. for radical expressions, how would you know if it's only one solution or two?

OpenStudy (anonymous):

@niiiice yes the common factor is 2.

OpenStudy (shamim):

welcome @niiice

OpenStudy (shamim):

i think it has only one solution

OpenStudy (shamim):

becausr 125/8 is a positive number

OpenStudy (anonymous):

ohh and if it's negative, there would be two solutions?

OpenStudy (shamim):

(-125/8)=(-5/2)*(-5/2)*(-5/2)

OpenStudy (shamim):

so if it is negative it has only one solution is -5/2

OpenStudy (anonymous):

@shamim how would you do 25x^2+5=14? I think I did mine wrong.. I'm really confused 25x^2+5=14 (5x+5)(5x-5)=14 5x+5=14 x=9/5 (5x-5)=14 x=19/5

OpenStudy (shamim):

yes its wrong

OpenStudy (anonymous):

@shamim what was my mistake?

OpenStudy (shamim):

u know a^2-b^2=(a+b)(a-b)

OpenStudy (shamim):

nd never a^2+b^2=(a+b)(a-b)

OpenStudy (anonymous):

@shamim okay? so do i need to square root 25x^2+5?

OpenStudy (shamim):